OpenStudy (midhun.madhu1987):

Use simplex method to solve the LPP: Max Z = 3x1 + 2x2 Subject to x1 + x2 <= 4 x1 - x2 <= 2 x1 , x2 >= 0 Can someone help me understand this step by step... ??

OpenStudy (anonymous):

iim interested in knowing what you study lol...

OpenStudy (midhun.madhu1987):

sure.. but somebody should be ready to explain :(

OpenStudy (anonymous):

sorry this is prolly beyond my scope...i dont know linear programming

OpenStudy (anonymous):

you should @Zarkon

OpenStudy (anonymous):

he knows everything....like everything

OpenStudy (midhun.madhu1987):

Thanks for that :)

OpenStudy (anonymous):

lol u didnt answer me

OpenStudy (midhun.madhu1987):

??

OpenStudy (anonymous):

what do you study?

OpenStudy (anonymous):

computer science? 3rd year?

OpenStudy (midhun.madhu1987):

ohh.. I am doing my Engineering in Computer Science..

OpenStudy (anonymous):

i knew it...

OpenStudy (midhun.madhu1987):

yep.. How you know that?

OpenStudy (anonymous):

the question...

OpenStudy (anonymous):

im in second year so lol....

OpenStudy (midhun.madhu1987):

coool

OpenStudy (anonymous):

@wio wanna take alook at this

OpenStudy (anonymous):

@wio accio

OpenStudy (anonymous):

harry potter magic doesnt work on him...

OpenStudy (perl):

this is a simple LPP Graph x1 + x2 <= 4 x1 - x2 <= 2 x1>=0 , x2 >= 0 (which means 1st quadrant)

OpenStudy (perl):

lets call x1 = x , and x2 = y x + y <= 4 x - y <= 2 graph y <= -x + 4 y >= x - 2

OpenStudy (perl):



OpenStudy (perl):